# Quantum Teleportation Coexisting with Classical Communications in Optical Fiber

### v3

@@
The ability for quantum and conventional networks to operate in the same optical fibers would aid the deployment of quantum network technology on a large scale. Quantum teleportation is a fundamental operation in quantum networking, but has yet to be demonstrated in fibers populated with high-power conventional optical signals. Here we report to the best of our knowledge the first demonstration of quantum teleportation over fibers carrying conventional telecommunications traffic. Quantum state transfer is achieved over a 30.2-km fiber carrying 400-Gbps C-band classical traffic with a Bell state measurement performed at the fiber’s midpoint. To protect quantum fidelity from spontaneous Raman scattering noise, we use optimal O-band quantum channels, narrow spectro-temporal filtering, and multi-photon coincidence detection. Fidelity is shown to be well maintained with an elevated C-band classical power of 18.7 dBm, which could support multiple classical channels totaling many terabits/s aggregate data rates. These results show the feasibility of advanced quantum and classical network applications operating within a unified fiber infrastructure.
@@ 1 Introduction
The study of quantum-classical coexistence has a long history with many experiments conducted for applications using weak coherent state (WCS) sources [3, 4, 5, 6, 7, 8, 9, 10, 11, 12], entangled photon pairs [13, 14, 15, 16, 17, 18, 19, 20, 21, 22], Bell state measurements (BSMs) on WCSs [23, 24], continuous variables [25], and squeezed light [26]. Although significant achievements and insights have been made in these studies, all experiments to date have focused on systems that directly transmit quantum information between network nodes. However, many next-generation quantum applications require the disembodied transfer of quantum states between users. Enabled by the non-local properties of quantum entanglement, quantum teleportation allows the transfer of a quantum state between two distant physical systems without the need for direct transmission [27]. It plays a foundational role in advanced applications such as quantum relays [28], quantum repeaters [29], networking quantum computers [30], and other applications in quantum science and technology [31].
@@
In this paper, we demonstrate a three-node quantum state teleportation system operating over 30.2 km of optical fiber that simultaneously carries high-power C-band classical communications at a rate of 400 Gbps. Teleportation is performed via a joint BSM on a single photon and one member of an entangled Bell state photon pair [40]. Utilizing wavelength multiplexing of quantum and classical signals, a heralded single photon carrying a polarization encoded qubit and one photon of an entangled Bell state photon pair are distributed over roughly 15 km of fiber to undergo a BSM near the midpoint of the fiber, whilst classical light traverses the full 30.2 km link plus an additional 48 km of deployed fiber.
We implement multiple techniques to mitigate degradation due to noise photons. At the quantum source nodes, we make use of the low SpRS noise in the O-band from strong C-band light [3], which has been shown successful for protecting quantum states from high-power C-band communications [9, 10, 11, 12, 24, 17]. More specifically, we allocate coexisting photons to the 1290-nm quantum channels to have nearly the lowest SpRS generation probability achievable using telecom-band wavelength engineering [17]. At quantum receivers, we use narrow-band spectro-temporal filtering to reject the detection of uncorrelated SpRS photons. Furthermore, each source uses spontaneous parametric down-conversion to generate photon pairs with strong temporal correlations to further reject noise via coincidence detection of the four photons. The design is also similar to a quantum relay, which has been predicted to reduce the impact of detector dark count noise [28] indicating a potential application in quantum-classical coexistence.
We first demonstrate that our approach allows high-power C-band communications to have no significant impact on entanglement distribution to the BSM node nor Hong-Ou-Mandel (HOM) interference [41] between independent heralded photons, which are the underlying operations of the quantum teleportation protocol [40]. We then perform quantum state teleportation and demonstrate that fidelity is maintained at elevated C-band launch powers as high as 74 mW (18.7 dBm), which is on the order of the highest achieved using direct transmission of quantum states [10, 11, 12, 17, 25]. Using a theoretical model, we then discuss the the pros and cons of using noisier quantum wavelength channels (eg., C-band), the effects of storage in imperfect quantum memories, and compare to methods using direct qubit transmission.
@@
Figure 1: (a) Conceptual diagram of the experiment. Alice prepares a qubit $|\psi\rangle_{A}$ that she wishes to transfer onto a photon at Bob’s node via quantum state teleportation. Alice encodes $|\psi\rangle_{A}$ onto a single photon, which is wavelength-division multiplexed into a fiber of length $L_{AC}$ to co-propagate with conventional classical communications traffic to a node at Charlie. The classical signal is de-multiplexed to bypass Charlie’s node and re-multiplexed to continue propagating over a fiber of length $L_{BC}$. Two photons generated at Bob’s node are prepared in an entangled Bell state, where one is transmitted to counter-propagate with the classical light to Charlie to undergo a Bell state measurement (BSM) with Alice’s qubit. After a BSM detection, Alice’s state is destroyed whilst the photon kept at Bob’s node is projected onto the state $\sigma|\psi\rangle_{A}$, where $\sigma$ is a unitary Pauli transformation unique to the BSM result. (b) Experimental implementation. Quantum and classical signals coexist in 30.2  km of fiber ($L_{AC}=15.2\text{ km}$, $L_{BC}=15.0\text{ km}$). Alice and Bob both generate photon pairs via non-degenerate spontaneous parametric down-conversion with center wavelengths of 1290 nm and 1310 nm. Alice’s source heralds a single photon, whereas Bob prepares a two-photon entangled Bell state. Qubits are encoded in polarization and the 1290-nm photons from each source are transmitted to Charlie for the BSM and teleportation is evaluated via four-fold coincidence detection and polarization analysis at Bob. The classical signal has a rate of 400-Gbps and is allocated to the 1547.32 nm C-band wavelength channel. This signal travels an additional 24-km of deployed fiber before and after multiplexing giving a total classical link distance of 78.2 km (FPC = fiber polarization controller, DWDM = dense wavelength division multiplexer, FWDM = O-band/C-band WDM, PPLN = periodically poled lithium niobate waveguide, $\lambda/2$ = half-wave plate, $\lambda/4$ = quarter-wave plate, LP = linear polarizer, LCR = liquid crystal retarder, PBS = polarizing beam splitter, FBS = 50:50 fiber coupler, VDL = variable optical delay line, FBG = fiber Bragg grating, CIR = circulator, $\text{D}_{j}$ = superconducting nanowire single photon detector, C = common port, P = pass port, R = reflect port, EDFA = erbium doped fiber amplifier).
@@ 2 Experimental Design
Both Alice and Bob generate their respective quantum signals via type-0 cascaded second harmonic generation-spontaneous parametric down conversion (c-SHG-SPDC) photon pair generation [42]. Alice and Bob each have independent periodically-poled lithium niobate (PPLN) waveguides that are phase matched for SHG at 650 nm. The cascaded second-order nonlinear process is analogous to four-wave mixing, which we pump using a 1300-nm continuous-wave distributed feedback laser that is intensity modulated using a lithium niobate Mach-Zehnder modulator to generate pulses with an approximate 65-ps temporal full-width at half maximum (FWHM) and a pulse repetition rate of 500 MHz. This pump is then split and directed to Alice and Bob’s nodes (see supp. materials). The 650-nm pulses generated inside each waveguide simultaneously drives SPDC to produce a two-mode squeezed vacuum state with a non-degenerate joint spectrum centered around 1300 nm. This state approximates a probabilistic source of single-photon pairs given the probability of generating a pair per pulse is $\ll 1$.
@@
To perform the BSM, we place PBSs in each output of the 50:50 coupler before the detectors $D_{1}$ and $D_{2}$ which are set to project onto orthogonal polarizations $|H\rangle_{D_{1}}$ and $|V\rangle_{D_{2}}$. A two-fold coincidence detection then indicates that Alice and Bob’s photons were in the $|\Psi^{-}\rangle_{AB}$ Bell state due to HOM interference effects on the other Bell states [40]. Due to the non-local correlations of Bob’s entangled photons in the state $|\Psi^{-}\rangle_{B}$, the destructive measurement of $|\Psi^{-}\rangle_{AB}$ results in the remaining photon at Bob’s node being projected onto Alice’s initial prepared state $|\psi\rangle_{A}$ [40]. Although teleportation can be performed without projecting onto the $H/V$ basis, Bell state analyzers that measure in the qubit’s eigenbasis can double the success probability of a BSM by also measuring the $|\Psi^{+}\rangle$ Bell state if detectors were added to the other ports of our PBSs [45]. Including the PBSs thus allows us to investigate BSM nodes most likely to be implemented in future networks. We also note that due to polarization mode dispersion over long fibers, the SpRS noise is largely unpolarized [46, 17] which means the polarizing elements roughly halve the amount of SpRS entering each detector.
@@
Since the classical data modulation is uncorrelated with the pulsed quantum sources in the time-domain, the coincidence detection window acts as a temporal filter of the SpRS noise. The temporal correlations in photon pair generation can also be exploited for time-domain filtering. Heralding single photons can improve tolerance to noise compared to sources such as WCSs by accepting a detection only when the heralding signal ensures a photon had been generated which reduces the probability of transmitting a vacuum state [47]. In our four-photon teleportation system, both Alice and Bob’s pair sources benefit from these correlations and midpoint BSM nodes in a quantum relay configuration have been predicted to offer higher tolerance to dark count effects [28]. Given this, we predict that instead of Alice directly transmitting her qubit to Bob, our system can potentially offer higher tolerance to classical power in quantum-classical networks (see supp. materials).
The coexistence study is conducted over spooled fiber in a laboratory located at Northwestern University’s Evanston campus. The transmission loss over the 30.2-km fiber is 4.9 dB at the classical signal’s 1547.32-nm wavelength and the 1290-nm quantum signals have a total loss of 10.1 dB (5.1 dB and 5.0 dB for Alice and Bob’s photons, respectively). At Charlie, the bypassing of the BSM by the classical signal using WDMs adds an additional 1.2 dB insertion loss. However, the classical transceiver is located at Northwestern’s Chicago campus which is connected to Evanston by a 24 km deployed fiber pair and thus operates over a total link distance of 30.2 + 48 km = 78.2 km. Due to the high loss over the full link (22.8 dB) and an initial output power of -9 dBm, amplification is needed to meet the classical system’s $-18$ dBm minimum received power requirements. We achieve this using erbium doped fiber amplifiers (EDFAs) placed just before multiplexing to co-propagate with Alice’s photon. The minimum power launched into the 30.2-km fiber to operate the 400-Gbps channel is $P_{\text{min}}=-3$ dBm (0.5 mW) for the receiver located 30.2 km + 24 km away from the EDFA, and -10.7 dBm (85 $\mu$W) if the classical system operated solely over the 30.2-km link. However, we explore elevated powers using the EDFAs to evaluate whether teleportation can coexist with more hostile classical systems such as ultra high-power long-distance DWDM conventional communications.
@@ 3.1 Noise rates, entanglement distribution, and Hong-Ou-Mandel interference
We now evaluate the performance of our quantum system. First, we characterize how much noise is introduced into the detectors $D_{1}$ and $D_{2}$ as a function of the classical power $P_{\text{cl}}$ that is multiplexed into the 30.2-km fiber. Fig. 2(a) shows the single photon count rates as we increase $P_{\text{cl}}$, where the SNSPD dark count rates have been subtracted. The noise rates scale linearly as 79.0 counts/s/mW and 97.9 counts/s/mW in $D_{1}$ and $D_{2}$, respectively, the difference is due to slightly unequal efficiencies after the 50:50 coupler. Since a single 400-Gbps channel can operate with as low as $P_{\text{cl}}=0.5$ mW, the system can operate with nearly no noticeable noise increase above the SNSPD dark count rate. Fig. 2(a) also includes an axis showing the ratio $M$ of $P_{\text{cl}}$ to the minimum single-channel power, which can be used to estimate the maximum aggregate data rates the $P_{\text{cl}}$ could support in a multi-channel system via $M\times 400$ Gpbs.
@@
Fig. 2(b) shows how Bob’s entanglement source’s non-local two-photon interference fringe visibility in the vertical and anti-diagonal bases is impacted after distributing one photon over 15.0 km to Charlie as we increase the C-band signal’s power launched into the full 30.2-km link. The visibility is defined as $V_{\rm ent}=(R_{\rm max}-R_{\rm min})/(R_{\rm max}+R_{\rm min})$, where $R_{\rm max(\rm min)}$ are the maximum (minimum) two-fold coincidence counts of the interference fringe. Fig. 2(b) shows no noticeable degradation of $V_{\rm ent}$ up to $P_{\text{cl}}=$ 74 mW, in which we measure $V_{\rm ent,V}=97.5\pm 0.1\%$ and $V_{\rm ent,A}=95.3\pm 0.2\%$. All error bars in this paper are determined via the Monte-Carlo method assuming Poisson photon counting statistics. These results indicate that entanglement distribution is well preserved and that visibility is mainly limited by multi-pair emission during SPDC. These values are also well above the $1/\sqrt{2}$ bound for demonstrating the non-local nature of quantum entanglement [50] and more than sufficient to use for quantum teleportation.
Next, we characterize HOM interference between heralded photons from Alice and Bob’s sources as we increase $P_{\text{cl}}$. To ensure polarization indistinguishably, we align the PBSs in the BSM setup to instead project onto $|H\rangle_{D_{1}}|H\rangle_{D_{2}}$ and prepare Alice and Bob’s heralded photons in the state $|H\rangle$. We then record four-fold coincidences as we vary the relative time-of-arrival of each photon using Alice’s variable optical delay. The quality of the interference is determined by the visibility of the fringe $V_{\rm HOM}=(R_{\rm max}-R_{\rm min})/R_{\rm max}$, where $R_{\rm max(min)}$ are the maximum (minimum) four-fold coincidence rates of the HOM interference fringe. The results are shown in Fig. 2(c). We measure a visibility of $V_{\rm HOM}=82.9\pm 4.5\%$ without the classical source, which is comparable to the $V_{\rm HOM}=83.1\pm 1.4\%$ visibility we obtain without distribution over the 30.2-km fiber (see supp. materials). We measure $V_{\rm HOM}=80.3\pm 3.8\%$ with $P_{\text{cl}}$ set to 74 mW, showing interference is not significantly degraded by high-power coexisting C-band light. Each visibility is well above the classical bound of 50% [51], indicating our ability to demonstrate non-classical interference across the range of the tested coexisting C-band powers.

### v4

@@
The ability for quantum and conventional networks to operate in the same optical fibers would aid the deployment of quantum network technology on a large scale. Quantum teleportation is a fundamental operation in quantum networking, but has yet to be demonstrated in fibers populated with high-power conventional optical signals. Here we report to the best of our knowledge the first demonstration of quantum teleportation over fibers carrying conventional telecommunications traffic. Quantum state transfer is achieved over a 30.2-km fiber carrying 400-Gbps C-band classical traffic with a Bell state measurement performed at the fiber’s midpoint. To protect quantum fidelity from spontaneous Raman scattering noise, we use optimal O-band quantum channels, narrow spectro-temporal filtering, and multi-photon coincidence detection. Fidelity is shown to be well maintained with an elevated C-band launch power of 18.7 dBm for the single-channel 400-Gbps signal, which we project could support multiple classical channels totaling many terabits/s aggregate data rates. These results show the feasibility of advanced quantum and classical network applications operating within a unified fiber infrastructure.
@@ 1 Introduction
The study of quantum-classical coexistence has a long history with many experiments conducted for applications using weak coherent state (WCS) sources [3, 4, 5, 6, 7, 8, 9, 10, 11, 12], entangled photon pairs [13, 14, 15, 16, 17, 18, 19, 20, 21, 22], Bell state measurements (BSMs) on WCSs [23, 24], continuous variables (CVs) [25], and squeezed light [26]. Although significant achievements and insights have been made in these studies, all experiments to date have focused on systems that directly transmit quantum information between network nodes. However, many next-generation quantum applications require the disembodied transfer of quantum states between users. Enabled by the non-local properties of quantum entanglement, quantum teleportation allows the transfer of a quantum state between two distant physical systems without the need for direct transmission [27]. It plays a foundational role in advanced applications such as quantum relays [28], quantum repeaters [29], networking quantum computers [30], and other applications in quantum science and technology [31].
@@
In this paper, we demonstrate a three-node quantum state teleportation system operating over 30.2 km of optical fiber that simultaneously carries high-power C-band classical communications at a rate of 400 Gbps. Teleportation is performed via a joint BSM on a single photon and one member of an entangled Bell state photon pair [40]. Utilizing wavelength multiplexing of quantum and classical signals, a heralded single photon carrying a polarization encoded qubit and one photon of an entangled Bell state photon pair are distributed over roughly 15 km of spooled fiber to undergo a BSM near the midpoint of the fiber, whilst classical light traverses the full 30.2-km link plus an additional 48 km of deployed fiber.
We implement multiple techniques to mitigate degradation due to noise photons. At the quantum source nodes, we make use of the low SpRS noise in the O-band from strong C-band light [3], which has been shown successful for protecting quantum states from high-power C-band communications [9, 10, 11, 12, 24, 17]. More specifically, we allocate coexisting photons to the 1290-nm quantum channels to have nearly the lowest SpRS generation probability achievable using telecom-band wavelength engineering [17]. At quantum receivers, we use narrow-band spectro-temporal filtering to reject the detection of uncorrelated SpRS photons. Furthermore, each source uses spontaneous parametric down-conversion to generate photon pairs with strong temporal correlations to further reject noise via coincidence detection of the four photons.
We first demonstrate that our approach allows high-power C-band communications to have no significant impact on entanglement distribution to the BSM node nor Hong-Ou-Mandel (HOM) interference [41] between independent heralded photons, which are the underlying operations of the quantum teleportation protocol [40]. We then perform quantum state teleportation and demonstrate that fidelity is maintained at elevated C-band launch powers as high as 74 mW (18.7 dBm), which is on the order of the highest achieved using direct transmission of quantum states [10, 11, 12, 17, 25]. Using a theoretical model, we then discuss the the pros and cons of using noisier quantum wavelength channels (e.g., C-band), the effects of storage in imperfect quantum memories, and compare to methods using direct qubit transmission.
@@
Figure 1: (a) Conceptual diagram of the experiment. Alice prepares a qubit $|\psi\rangle_{A}$ that she wishes to transfer onto a photon at Bob’s node via quantum state teleportation. Alice encodes $|\psi\rangle_{A}$ onto a single photon, which is wavelength-division multiplexed into a fiber of length $L_{AC}$ to co-propagate with conventional classical communications traffic to a node at Charlie. The classical signal is de-multiplexed to bypass Charlie’s node and re-multiplexed to continue propagating over a fiber of length $L_{BC}$. Two photons generated at Bob’s node are prepared in an entangled Bell state, where one is transmitted to counter-propagate with the classical light to Charlie to undergo a Bell state measurement (BSM) with Alice’s qubit. After a BSM detection, Alice’s state is destroyed whilst the photon kept at Bob’s node is projected onto the state $\sigma|\psi\rangle_{A}$, where $\sigma$ is a unitary Pauli transformation identifiable from the BSM result. (b) Experimental implementation. Quantum and classical signals coexist in 30.2  km of spooled fiber ($L_{AC}=15.2\text{ km}$, $L_{BC}=15.0\text{ km}$). Alice and Bob both generate photon pairs via non-degenerate spontaneous parametric down-conversion with center wavelengths of 1290 nm and 1310 nm. Alice’s source heralds a single photon, whereas Bob prepares a two-photon entangled Bell state. Qubits are encoded in polarization and the 1290-nm photons from each source are transmitted to Charlie for the BSM. Teleportation is evaluated via four-fold coincidence detection and polarization analysis of Bob’s target photon. The classical source transmits a 400-Gbps C-band signal (1547.32 nm) over 24 km of deployed fiber before being amplified and multiplexed into the 30.2-km spool. After de-multiplexing this propagates over another 24-km deployed fiber to the receiver, totaling 78.2 km of fiber. (FPC = fiber polarization controller, DWDM = dense wavelength division multiplexer, FWDM = O-band/C-band WDM, PPLN = periodically poled lithium niobate waveguide, $\lambda/2$ = half-wave plate, $\lambda/4$ = quarter-wave plate, LP = linear polarizer, LCR = liquid crystal retarder, PBS = polarizing beam splitter, FBS = 50:50 fiber coupler, VDL = variable optical delay line, FBG = fiber Bragg grating, CIR = circulator, $\text{D}_{j}$ = superconducting nanowire single photon detector, C = common port, P = pass port, R = reflect port, EDFA = erbium doped fiber amplifier).
@@ 2 Experimental design
Both Alice and Bob generate their respective quantum signals via type-0 cascaded second harmonic generation-spontaneous parametric down conversion (c-SHG-SPDC) photon pair generation [42]. Alice and Bob each have independent periodically-poled lithium niobate (PPLN) waveguides that are phase matched for SHG at 650 nm. The cascaded second-order nonlinear process is analogous to four-wave mixing, which we pump using a 1300-nm continuous-wave distributed feedback laser that is intensity modulated using a lithium niobate Mach-Zehnder modulator to generate pulses with an approximate 65-ps temporal full-width at half maximum (FWHM) and a pulse repetition rate of 500 MHz. This pump is then split and directed to Alice and Bob’s nodes (see supp. materials). The 650-nm pulses generated inside each waveguide simultaneously drives SPDC to produce a superposition of photon pairs with a non-degenerate spectrum centered around 1300 nm and forms the basis for Alice’s heralded single-photon source (HSPS) and generating Bob’s entangled photons, but requires a low single-photon pair generation rate to minimize the probability of multi-photon emission [43]
@@
To perform the BSM, we place PBSs in each output of the 50:50 coupler before the detectors $D_{1}$ and $D_{2}$ which are set to project onto orthogonal polarizations $|H\rangle_{D_{1}}$ and $|V\rangle_{D_{2}}$. A two-fold coincidence detection then indicates that Alice and Bob’s photons were in the $|\Psi^{-}\rangle_{AB}$ Bell state due to HOM interference effects on the other Bell states [40]. Due to the non-local correlations of Bob’s entangled photons in the state $|\Psi^{-}\rangle_{B}$, the destructive measurement of $|\Psi^{-}\rangle_{AB}$ results in the remaining photon at Bob’s node being projected onto Alice’s initial prepared state $|\psi\rangle_{A}$ [40]. Although teleportation can be performed without projecting onto the $H/V$ basis, Bell state analyzers that measure in the qubit’s eigenbasis can double the success probability of a BSM by also measuring the $|\Psi^{+}\rangle$ Bell state if detectors were added to the other ports of our PBSs [46]. Including the PBSs thus allows us to investigate BSM nodes most likely to be implemented in future networks. We note that SpRS photons in long-distance fibers are often unpolarized due to polarization mode dispersion [47, 17] meaning polarizing elements roughly halve the noise rates. This was confirmed in our system by polarization analysis of the noise photons at the BSM node (see supp. materials).
@@
Since the classical data modulation is uncorrelated with the pulsed quantum sources in the time-domain, the coincidence detection window acts as a temporal filter of the SpRS noise. The temporal and photon-number correlations of our photon-pair sources further help the system’s tolerance to noise by accepting detection events at the receiver only in the time slots when a detection in the heralding arm indicated a photon was transmitted over the noisy channel [48]. This can significantly improve the signal-to-noise ratio compared to non-heralded sources that have a high probability of transmitting a vacuum state such as WCSs. In our four-photon teleportation system, both Alice and Bob’s pair sources benefit from these correlations.
The coexistence study is conducted over spooled fiber in a laboratory located at Northwestern University’s Evanston campus. The transmission loss over the 30.2-km fiber is 4.9 dB at the classical signal’s 1547.32-nm wavelength and the 1290-nm quantum signals have a total loss of 10.1 dB (5.1 dB and 5.0 dB for Alice and Bob’s photons, respectively). At Charlie, the bypassing of the BSM by the classical signal using WDMs adds an additional 1.2 dB insertion loss. However, the classical transceiver is located at Northwestern’s Chicago campus which is connected to Evanston by a 24-km deployed fiber pair and thus operates over a total link distance of 30.2 + 48 km = 78.2 km. Due to the high loss over the full link (22.8 dB) and an initial output power of -9 dBm, amplification is needed to meet the classical system’s $-18$ dBm minimum received power requirements. We achieve this using erbium doped fiber amplifiers (EDFAs) placed just before multiplexing to co-propagate with Alice’s photon. The minimum power launched into the 30.2-km fiber to operate the 400-Gbps channel is $P_{\text{min}}=-3$ dBm (0.5 mW) for the receiver located 30.2 km + 24 km away from the EDFA, and -10.7 dBm (85 $\mu$W) if the classical system operated solely over the 30.2-km link. However, we explore elevated powers using the EDFAs to evaluate whether teleportation can coexist with more hostile classical systems such as ultra high-power long-distance DWDM conventional communications.
@@ 3.1 Noise rates, entanglement distribution, and Hong-Ou-Mandel interference
We now evaluate the performance of our quantum system. First, we characterize how much noise is introduced into the detectors $D_{1}$ and $D_{2}$ as a function of the classical power $P_{\text{cl}}$ that is multiplexed into the 30.2-km fiber. Fig. 2(a) shows the single photon count rates as we increase $P_{\text{cl}}$, where the SNSPD dark count rates have been subtracted. The noise rates scale linearly as 79.0 counts/s/mW and 97.9 counts/s/mW in $D_{1}$ and $D_{2}$, respectively, the difference is due to slightly unequal efficiencies after the 50:50 coupler. Since a single 400-Gbps channel can operate with as low as $P_{\text{cl}}=0.5$ mW, the system can operate with nearly no noticeable noise increase above the SNSPD dark count rate. Fig. 2(a) also shows the ratio $M=P_{\text{cl}}/P_{\text{min}}$, which is often used to roughly estimate the maximum number of channels $M$ that $P_{\text{cl}}$ could support across multiple WDM channels (e.g., [23, 24]) assuming each operates at $P_{\text{min}}$ and has similar SpRS properties (see supp. materials).
@@
Fig. 2(b) shows how Bob’s entanglement source’s non-local two-photon interference fringe visibility in the vertical and anti-diagonal bases is impacted after distributing one photon over 15.0 km to Charlie as we increase the C-band signal’s power launched into the full 30.2-km link. The visibility is defined as $V_{\rm ent}=(R_{\rm max}-R_{\rm min})/(R_{\rm max}+R_{\rm min})$, where $R_{\rm max(\rm min)}$ are the maximum (minimum) two-fold coincidence counts of the interference fringe. Fig. 2(b) shows no noticeable degradation of $V_{\rm ent}$ up to $P_{\text{cl}}=$ 74 mW, in which we measure $V_{\rm ent,V}=97.5\pm 0.1\%$ and $V_{\rm ent,A}=95.3\pm 0.2\%$. All values are determined by raw data without accidental count subtraction and error bars are calculated via the Monte-Carlo method assuming Poisson photon counting statistics. These results indicate that entanglement distribution is well preserved and that visibility is mainly limited by multi-pair emission during SPDC. These values are also well above the $1/\sqrt{2}$ bound for demonstrating the non-local nature of quantum entanglement [50] and more than sufficient to use for quantum teleportation.
Next, we characterize HOM interference between heralded photons from Alice and Bob’s sources as we increase $P_{\text{cl}}$. To ensure polarization indistinguishably, we align the PBSs in the BSM setup to instead project onto $|H\rangle_{D_{1}}|H\rangle_{D_{2}}$ and prepare Alice and Bob’s heralded photons in the state $|H\rangle$. We then record four-fold coincidences as we vary the relative time-of-arrival of each photon using Alice’s variable optical delay. The quality of the interference is determined by the visibility of the fringe $V_{\rm HOM}=(R_{\rm max}-R_{\rm min})/R_{\rm max}$, where $R_{\rm max(min)}$ are the maximum (minimum) four-fold coincidence rates of the HOM interference fringe. The results are shown in Fig. 2(c). We measure a visibility of $V_{\rm HOM}=82.9\pm 4.5\%$ without the classical source, which is comparable to the $V_{\rm HOM}=83.1\pm 1.4\%$ visibility we obtain without distribution over the 30.2-km fiber (see supp. materials). We measure $V_{\rm HOM}=80.3\pm 3.8\%$ with $P_{\text{cl}}$ set to 74 mW, which shows that interference is well preserved in the presence of high-power C-band light but is still limited by the imperfect spectral and single-photon purity of our sources. Each visibility is well above the classical bound of 50% [51], indicating our ability to demonstrate non-classical interference across the range of the tested coexisting C-band powers.